# A quantum information theoretic analysis of reinforcement learning-assisted quantum architecture search

### v1

@@ III.1.2 Anti-correlation between entanglement bounds
We present in Figure 2, the variation of $\eta_{0}^{k}$ (in blue) and $\eta_{1}^{k}$ (in red) for different values of $k$. We observe that $\eta_{1}^{k}\approx-1~{}\forall k\in(i,j)$. This indicates strong cumulative anti-correlation between the upper and lower bounds of concurrence. For $\eta_{0}^{k}$, we observe a gradual change of $\text{PCC}^{kj}_{ul}$ from -0.582 to 1. We note that $\eta_{0}^{k}$ changes from $-0.0451$ to $0.279$ for $k\in[0.313,0.379]$ with $\eta_{0}^{k}=0$ at $k=k_{\ast}=0.322$. We call this point the phase transition point of correlation between upper and lower bounds of concurrence for the VQSD ansatz. At values of $k<k_{\ast}$, there is anti-correlation between the upper and lower bounds of concurrence of the ansatz for both the intervals $\mathbbm{u}_{1}$ and $\mathbbm{u}_{2}$. While for values of $k>k_{\ast}$, there is (anti-)correlation between the upper and lower bounds of concurrence of the ansatz for both the intervals $(\mathbbm{u}_{2})~{}\mathbbm{u}_{1}$.
III.1.3 Agent and environment interaction reduction based on entanglement bound
Figure 3: We observe that the RL-agent can generate ansatz with very high concurrence with less number of 2-qubit gates and circuit depth compared to the ansatzes with very low concurrence. As the main aim of RL-VQSD is to diagonalise unitary with a small number of gates and depth, our analysis shows that if we particularly focus on admissible ansatzes near the upper bound of concurrence, we can not only reduce the computational cost of RL-VQSD severely but find optimal admissible ansatzes. The numbers in the bars correspond to the amount of entanglement generated by the RL-ansatz when initialised in the vacuum state.
Our investigation reveals that the admissible ansatzes corresponding to the upper bound of concurrence has, on average, a shorter depth of $13.23$ and a smaller number of gates of $18.5$ compared to the admissible ansatzes corresponding to the lower bound of concurrence where the depth is $13.9$ and number of gates is $19$ over $25$ random quantum states. Moreover, Figure 3 makes it more prominent that the minimum number of gates and the depth of the admissible ansatzes are smaller for the upper bound of concurrence than the lower. This observation is crucial in reducing the subspace of all possible admissible ansatzes. The main motivation of RL-QAS in VQA is to find ansatzes that solve a VQA with a minimum number of gates and circuit depth; in the case of RL-VQSD, we can narrow down the search for an optimal admissible ansatz near the upper bound of entanglement by encouraging the RL-anent to produce ansatzes with very high concurrence and penalise for the other structures of ansatz.
@@ III.2.1 Contribution of individual qubits for change in entanglement
Figure 4: Relative change in the contribution of qubits 0 and 1 as a function of change in entanglement. Weight of the positive part $5.524$ and that of the negative part is $-4.854$.
@@
We present the variation of $\Delta_{q_{0}}^{q_{1}}$ for the change in concurrence due to the ansatz given by $\Delta_{c}\coloneqq C^{\rho_{q_{0}q_{1}}}-C^{\rho^{\prime}_{q_{0}q_{1}}}$ in Figure 4. We observe that the distribution of $\Delta_{q_{0}}^{q_{1}}$ is bounded in the interval $[-0.7,0.7]$ and is uniform with respect to $\Delta_{q_{0}}^{q_{1}}=0$. This implies an equal contribution to the VQSD ansatz from both the qubits for the range of change in entanglement. In the region of $\Delta_{q_{0}}^{q_{1}}>0$, the relative contribution of $q_{0}$ first increases and then decreases. However, no such pattern is observed for the region of $\Delta_{q_{0}}^{q_{1}}<0$. When $\Delta_{c}>0.4$, the value of $\Delta_{q_{0}}^{q_{1}}$ monotonically decreases with increase in $\Delta_{c}$. Also, for $\Delta_{c}<0.2$, $\Delta_{q_{0}}^{q_{1}}$ increases monotonically with increase in $\Delta_{c}$.
III.2.2 Problem with variational quantum state diagonalization
@@
Figure 5: The conditional entropy correlation between the first and the second qubit decreases as the magnitude of the largest and the 2nd largest eigenvalues increases. Meanwhile, the same correlation between the qubits increases with the magnitude of the 3rd and the smallest eigenvalues. The RL-ansatz proposes specific structures of admissible ansatzes, and for each such ansatz, it is not possible to have two different kinds of correlations. Hence, our investigation suggests that it is not feasible to find the smallest eigenvalues with the same ansatz that provides a good approximation of the largest eigenvalues. The correlation of conditional entropy between qubits is calculated using Pearson correlation function (PCC) and is the median over all the configuration of RL-ansatz proposed by the agent. The $\bigstar$ signifies the eigenvalues corresponding to the same state.
@@
Figure 6: The correlation in conditional entropy among qubits is in the range of mild correlation for $21$ random quantum states. This observation, along with the observation in Figure 5 certify that RL-agent only proposes a particular configuration of admissible ansatz that can find only the largest eigenvalues. The average correlation among all the $21$ states is $-0.4263$, which is in the regime of mild anti-correlation.
@@ IV Discussion
We observe that the concurrence of the admissible ansatz ranges between an upper and lower bound, which are independent of the initialisation of the weights of the deep-Q network. The upper and lower bounds are initially anti-correlated with respect to the initial state entanglement. The anti-correlation turns into a mild and eventually strong correlation as the entanglement of the initial state surpasses beyond a phase change point at concurrence $0.322$. We also see that the optimal configuration of the admissible ansatz lies in the upper bound of concurrence, which has a smaller circuit depth and requires fewer gates than the lower bound. These observations provide insight regarding the relation between the entanglement of the initial state and the entanglement generated by the RL-VQSD admissible ansatzes. We effectively utilise these observations to greatly reduce the RL-agent and environment interaction, i.e., the computational time needed to solve the RL-VQSD problem. Additionally, the admissible ansatzes not only diagonalize the random quantum states with high accuracy but can be used to generate close to maximum entangled states from the vacuum state.
@@
For future work, it would be interesting to see the implication of introducing entanglement Horodecki et al. (2009) or nonlocality Brunner et al. (2014); Sadhu and Das (2023)-based reward functions of the RL-agent for solving the VQSD problem. Specifically, given that most of the admissible ansatzes have high values of concurrence, it is expected that an entanglement or non-locality-based reward function may further reduce the time required to find the optimal configurations. Also, we may expect this to provide an optimal quantum circuit that provides both the lowest and highest eigenvalues of the input quantum state.
@@ Acknowledgements.
A. Sadhu acknowledges the PhD fellowship from the Raman Research Institute, Bangalore. AK would like to thank Mateusz Ostaszewski for the valuable discussions. This research was partially supported by PL-Grid Infrastructure Grant Number PLG/2023/016205.
@@ A.3 Correlation coefficient
$\textrm{PCC}_{r_{1}r_{2}}=\frac{\mathrm{E}[(r_{1}-\mu_{r_{1}})(r_{2}-\mu_{r_{2}})]}{\sigma_{r_{1}}\sigma_{r_{2}}},$ (9)
where $\mu_{r_{j}}$, $\sigma_{r_{j}}$ are the mean and standard deviation of the distribution $r_{j}$ with $j\in\{1,2\}$ and $\mathrm{E}[.]$ is an operator that returns the expectation value of a random variable. Note that the Pearson correlation coefficient (PCC) is symmetric: $\textrm{PCC}_{r_{1}r_{2}}=\textrm{PCC}_{r_{2}r_{1}}$. The value of PCC ranges in between $-1$ to $+1$. A correlation value of +1(1)11+1(-1)+ 1 ( - 1 ) implies perfect correlation (anti-correlation) between the two random variables $r_{1}$ and $r_{2}$, while zero implies no linear dependence between the variables.
@@ Appendix B The reinforcement learning enhanced variational quantum state diagonalisation
where $\rho_{\text{diag}}$ is $\rho$ diagonalized in its eigenbasis and $\vec{\theta}_{\text{opt}}$ are the optimal angles. Classical gradient-based methods such as SPSA and Gradient Descent, or gradient-free optimisation Powell (1994, 2006), can be utilized in the training process. (2) In the eigenvalue readout subroutine, using the optimised unitary $U(\vec{\theta}_{\text{opt}})$ and one copy of state $\rho$, one can extract all the eigenvalues for low-rank states or the largest eigenvalues for full-rank states. This is achieved by measuring $\rho^{\prime}$ in the computational basis, $\mathbf{b}=b_{1}b_{2}\ldots b_{n}$, as follows
@@ Appendix C Analysis of the Haar random input quantum states
Figure 7: The difference in conditional entropy converges towards the zero as the entanglement of the sampled input state increases. Meanwhile, as the input state gets closer to maximally entangled state (with concurrence $1$), the magnitude of the largest eigenvalue increases while the other eigenvalue decreases.
We see that the difference in conditional entropy in qubits of the input state varies evenly on either side of $S_{AB}^{\rho_{AB}}=0$ where $A$ and $B$ are subsystems containing one qubit each and $\rho_{AB}$ is the sampled Haar random state. Meanwhile, the area of the distribution of $S_{AB}^{\rho_{AB}}$ on the ether side of zero entropy converges towards zero as the input state entanglement increases. Furthermore, the magnitude of the input quantum state’s largest eigenvalue increases with the increase in entanglement, and all the other eigenvalues converge towards 00.

### v2

@@ III.1.3 Enhancing the performance of agent based on initial state entanglement
where the first, Entanglement Enhancing (EE) block, helps the diagonalising state to achieve the highest entanglement before it is fed to the second part, $U_{\text{RL-agent}}$, which is decided by the RL-agent. For example, for seed $27$ (see Appendix C for details), we get a random quantum state with concurrence $0.198$, and using the EE, we can increase the input state concurrence to $0.215$ at $\theta=0.5$ as shown in Fig. 4.
Figure 4: The concurrence of input state with $\theta$ of CRX($\theta$) gate in Entanglement Enhancing increases the concurrence of the input random state from $0.198$ to $0.215$ when $\theta=0.5$.
Our investigation reveals that turning on the EE block with $\theta=0.5$ enhances the performance of the RL-agent by $2$-times compared to the default setting at $\theta=0$. The results are summarised in the Tab. 2 where we see that the number of successful episodes and the reward acquired (where the reward function is in the form defined in Eq. 12) by the agent in the RL-VQSD process increase by $2$-fold while the EE block is activated for $\theta=0.5$ compared to default settings at $\theta=0$.
Table 2: Turning on the Entanglement Enhancing (EE) mode (denoted as Enhanced ($\theta=0.5$)) enhances the learning of the performance of the RL-agent in the absence of the EE mode (denoted as Default ($\theta=0$)). In the first $3500$, the Enhanced ($\theta=0.5$) collects $2.08$ times more reward (tot. rwd. coll.) and $2.22$ times more successful episodes (Tot. succ. ep.) than Default ($\theta=0$). Moreover, the average number of 2-qubit (2q gate) and depth over the first $100$ successful episodes reduced in the case of Enhanced ($\theta=0.5$) settings compared to Default ($\theta=0$), showing enhanced learning of the RL-agent for entanglement enhanced input state.
EE mode Tot. succ. Tot. rwd. 1st 100 succ. ep. (avg.)
($\theta$) ep. coll. 1q gate 2q gate depth
Enhanced ($\theta=0.5$) $858$ $4.45\times 10^{5}$ $15.41$ $4.02$ $14.6$
Default ($\theta=0$) $387$ $2.14\times 10^{5}$ $14.71$ $5.29$ $15.59$
Meanwhile, in the Tab. 3, we show that the EE block also helps to optimise the cost function and the minimum number of 2-qubit gates in the optimal admissible ansatz compared to the default settings.
Table 3: Turning on the Entanglement Enhancing (EE) mode (denoted as Enhanced ($\theta=0.5$)) optimises the number of minimum number of 2-qubit gates (Min. 2q gate) in RL-ansatz better than in the absence of the EE mode (denoted as Default ($\theta=0$)). Moreover, the Enhanced ($\theta=0.5$) mode minimises the cost function better than the Default ($\theta=0$) mode.
EE mode ($\theta$) Cost function Min. 1q gate Min. 2q gate
Enhanced ($\theta=0.5$) $5.12\times 10^{-7}$ $14$ $2$
Default ($\theta=0$) $6.08\times 10^{-7}$ $11$ $5$
The aforementioned observations lead us to the conclusion that maximising the entanglement of the input state through the initialisation strategy of Fig. 2 enhances the learning process of the RL-agent and facilitates the further optimisation of the admissible ansatzes compared to the start-of-art Kundu (2024).
@@ III.2.1 Contribution of individual qubits for change in entanglement
Figure 5: Relative change in the contribution of qubits 0 and 1 as a function of change in entanglement. We quantify the relative contribution of individual qubits to the RL-VQSD ansatz using $\Delta_{q_{0}}^{q_{1}}$ (see Eq. (3)). For the change in concurrence due to the ansatz, the distribution of $\Delta_{q_{0}}^{q_{1}}$ is uniform with respect to $\Delta_{q_{0}}^{q_{1}}=0$. The magnitude of the cumulative weight of all data points in the region $\Delta_{q_{0}}^{q_{1}}<0$ is $4.854$ while that for data points in the region $\Delta_{q_{0}}^{q_{1}}>0$ is $5.524$. This indicates equal contribution from both the qubits to the RL-VQSD ansatz for the range of change in entanglement.
@@
We present the variation of $\Delta_{q_{0}}^{q_{1}}$ for the change in concurrence due to the ansatz given by $\Delta_{c}\coloneqq C^{\rho_{q_{0}q_{1}}}-C^{\rho^{\prime}_{q_{0}q_{1}}}$ in Figure 5. We observe that the distribution of $\Delta_{q_{0}}^{q_{1}}$ is bounded in the interval $[-0.7,0.7]$ and is uniform with respect to $\Delta_{q_{0}}^{q_{1}}=0$. The magnitude of the cumulative weight of all data points in the region $\Delta_{q_{0}}^{q_{1}}<0$ is $4.854$ while that for data points in the region $\Delta_{q_{0}}^{q_{1}}>0$ is $5.524$. This indicates equal contribution from both the qubits to the RL-VQSD ansatz for the range of change in entanglement. In the region of $\Delta_{q_{0}}^{q_{1}}>0$, the relative contribution of $q_{0}$ first increases and then decreases. However, no such pattern is observed for the region of $\Delta_{q_{0}}^{q_{1}}<0$. When $\Delta_{c}>0.4$, the value of $\Delta_{q_{0}}^{q_{1}}$ monotonically decreases with increase in $\Delta_{c}$. Also, for $\Delta_{c}<0.2$, $\Delta_{q_{0}}^{q_{1}}$ increases monotonically with increase in $\Delta_{c}$.
III.2.2 Problem with variational quantum state diagonalisation
@@
Figure 6: The conditional entropy correlation between the first and the second qubit decreases as the magnitude of the largest and the second largest eigenvalues increases. Meanwhile, the same correlation between the qubits increases with the magnitude of the third and the smallest eigenvalues. The RL-ansatz proposes specific structures of admissible ansatzes, and for each such ansatz, it is not possible to have two different kinds of correlations. Hence, our investigation suggests that it is not feasible to find the smallest eigenvalues with the same ansatz that provides a good approximation of the largest eigenvalues. The correlation of conditional entropy between qubits is calculated using Pearson correlation coefficient (PCC) and is the median over all the configuration of RL-ansatz proposed by the agent. The $\bigstar$ signifies the eigenvalues corresponding to the same state.
@@
Figure 7: The correlation in conditional entropy among qubits is in the range of mild correlation for $21$ random quantum states. This observation, along with the observation in Figure 6 certify that RL-agent only proposes a particular configuration of admissible ansatz that can find only the largest eigenvalues. The average correlation among all the $21$ states is $-0.4263$ (horizontal black line), which is in the regime of mild anti-correlation.
@@ IV Discussion
We observe that the concurrence of the admissible ansatz ranges between an upper and lower bound, which are independent of the initialisation of the weights of the deep-Q network. The upper and lower bounds are initially anti-correlated with respect to the initial state entanglement. The anti-correlation turns into a mild and eventually strong correlation as the entanglement of the initial state surpasses beyond a phase change point at concurrence $0.322$. We also see that the optimal configuration of the admissible ansatz lies in the upper bound of concurrence, which has a smaller circuit depth and requires fewer gates than the lower bound. These observations provide insight regarding the relation between the entanglement of the initial state and the entanglement generated by the RL-VQSD admissible ansatzes. We effectively utilise these observations to greatly reduce the RL-agent and environment interaction, i.e., the computational time needed to solve the RL-VQSD problem. Specifically, we propose an entanglement enhancing initial block that adjusts the input state’s concurrence thereby enhancing the performance of the RL-agent by a factor of $2$ Additionally, the admissible ansatzes not only diagonalise the random quantum states with high accuracy but can be used to generate close to maximum entangled states from the vacuum state.
@@
For future work, it would be interesting to conduct more comprehensive study of the implication of introducing entanglement Horodecki et al. (2009) or nonlocality Brunner et al. (2014); Sadhu and Das (2023)-based reward functions of the RL-agent in QAS. Specifically, given that most of the admissible ansatzes have high values of concurrence, it is expected that an entanglement or non-locality-based reward function may further reduce the time required to find the optimal configurations. Also, we may expect this to provide an optimal quantum circuit that provides both the lowest and highest eigenvalues of the input quantum state.
@@ Acknowledgements.
A. Sadhu acknowledges the Ph.D. fellowship from the Raman Research Institute, Bangalore. AK would like to thank Jarosław Miszczak and Mateusz Ostaszewski for their valuable comments on the manuscript and discussions. This research was partially supported by PL-Grid Infrastructure Grant Number PLG/2023/016205.
@@ A.3 Correlation coefficient
$\textrm{PCC}^{r_{1}r_{2}}=\frac{\mathrm{E}[(r_{1}-\mu_{r_{1}})(r_{2}-\mu_{r_{2}})]}{\sigma_{r_{1}}\sigma_{r_{2}}},$ (10)
where $\mu_{r_{j}}$, $\sigma_{r_{j}}$ are the mean and standard deviation of the distribution $r_{j}$ with $j\in\{1,2\}$ and $\mathrm{E}[.]$ is an operator that returns the expectation value of a random variable. Note that the Pearson correlation coefficient (PCC) is symmetric: $\textrm{PCC}^{r_{1}r_{2}}=\textrm{PCC}^{r_{2}r_{1}}$. The value of PCC ranges in between $-1$ to $+1$. A correlation value of +1(1)11+1(-1)+ 1 ( - 1 ) implies perfect correlation (anti-correlation) between the two random variables $r_{1}$ and $r_{2}$, while zero implies no linear dependence between the variables.
@@ Appendix B The reinforcement learning enhanced variational quantum state diagonalisation
where $\rho_{\text{diag}}$ is $\rho$ diagonalised in its eigenbasis and $\vec{\theta}_{\text{opt}}$ are the optimal angles. Classical gradient-based methods such as SPSA and Gradient Descent, or gradient-free optimisation Powell (1994, 2006), can be utilised in the training process. (2) In the eigenvalue readout subroutine, using the optimised unitary $U(\vec{\theta}_{\text{opt}})$ and one copy of state $\rho$, one can extract all the eigenvalues for low-rank states or the largest eigenvalues for full-rank states. This is achieved by measuring $\rho^{\prime}$ in the computational basis, $\mathbf{b}=b_{1}b_{2}\ldots b_{n}$, as follows
@@ Appendix C Analysis of the Haar random input quantum states
Figure 8: The difference in conditional entropy converges towards the zero as the entanglement of the sampled input state increases. Meanwhile, as the input state gets closer to the maximally entangled state (with concurrence $1$), the magnitude of the largest eigenvalue increases while the other eigenvalue decreases.
In Fig. 8, see that the difference in conditional entropy in qubits of the input state varies evenly on either side of $S_{AB}^{\rho_{AB}}=0$ where $A$ and $B$ are subsystems containing one qubit each and $\rho_{AB}$ is the sampled Haar random state. Meanwhile, the area of the distribution of $S_{AB}^{\rho_{AB}}$ on the ether side of zero entropy converges towards zero as the input state entanglement increases. Furthermore, the magnitude of the input quantum state’s largest eigenvalue increases with the increase in entanglement, and all the other eigenvalues converge towards 00. It also shows that even for $100000$ seeds, the probability of sampling states with concurrence greater than $0.6$ is very low, justifying the low density of points after $0.6$ in Fig. 1.
Sampling a random quantum state
If $n$ is the number of qubits and $s$ is the random seed to sample, then we can sample a state from random_density_matrix of IBM qiskit’s quantum_info module (in python) as follows
[…]
The state used in III.1, is sampled using $n=2$ and $s=27$